# Efficient Segmentation of Search Markets

### v2

@@ 4. $(iv)$
where the first equality follows from the definition of $\sigma^{PS}$ and the characterization of $\tau^{FB}_{\sigma^{PS}}$ in Proposition 1, the second equality follows by the assumption on $\lambda$ given in Proposition 4 and the fact that $f$ is the inverse of the mapping $t\mapsto 1/m^{\prime}(t)$, and the last equality follows by the construction of $\eta_{\sigma^{PS}}$ as given in Proposition 1. Thus, the pair $(\tau^{FB}_{\sigma^{PS}},\lambda(\frac{\eta_{\sigma^{PS}}}{\beta})\eta_{\sigma^{PS}})$ also satisfies (4) of Definition 1.
We can therefore conclude that $(\tau^{FB}_{\sigma^{PS}},\lambda(\frac{\eta_{\sigma^{PS}}}{\beta})\eta_{\sigma^{PS}})$ is a search equilibrium of $\sigma^{PS}$. However, $(\tau^{*}_{\sigma^{PS}},u^{*}_{\sigma^{PS}})$ is the essentially-unique search equilibrium of $\sigma^{PS}$. Thus, $\tau^{*}_{\sigma^{PS}}=\tau^{FB}_{\sigma^{PS}}$ for $\sigma_{\mu}^{PS}$-almost everywhere, and $u^{*}_{\sigma^{PS}}=\lambda(\frac{\eta_{\sigma^{PS}}}{\beta})\eta_{\sigma^{PS}}$. Since $(\sigma^{PS},\tau^{FB}_{\sigma^{PS}})$ solves (FB), the almost-everywhere equivalence between $\tau^{*}_{\sigma^{PS}}$ and $\tau^{FB}_{\sigma^{PS}}$ implies that $(\sigma^{PS},\tau^{*}_{\sigma^{PS}})$ also solves (FB), giving us the desired result. ∎
@@
[…]
Let ${V}(u)$ be the value function corresponding to (LP-$u$), and let $\mathcal{H}(u)$ be the set of maximizers. Since the objective is a continuous function and $\operatorname*{MPC}(F)$ is compact, $V(u)$ is continuous in $u$, and $\mathcal{H}(u)$ is non-empty and compact-valued for each $u>0$. Importantly, by construction, $H$ satisfies (5) if and only if $H\in\mathcal{H}(u_{H}^{*})$.
@@
[…]
@@
First, suppose the mapping $t\mapsto t/m(t)$ is concave, and therefore, $g$ is convex. In this case, $\phi(x,u)=g(\frac{\ell\cdot x}{u})\mathbbm{1}_{[\beta\ell x>u]}$ is convex in $x$ for all $u\in(0,\beta\ell)$. [^12] Recall that $\lim_{y\to 1/\beta}g(y)=0$. As a result, for all $H^{\prime},H^{\prime\prime}\in\operatorname*{MPC}(F)$ where $H^{\prime}$ is a mean-preserving contraction of $H^{\prime\prime}$, we have
[…]
@@
Second, suppose the mapping $t\mapsto t/m(t)$ is convex, and therefore, $g$ is concave. In this case, $\phi(x,u)=g(\frac{\ell\cdot x}{u})\mathbbm{1}_{[\ell\beta x>u]}$ is piece-wise concave in $x$ for all $u\in(0,\beta\ell)$. Specifically, it is constant over $[0,u/(\beta\ell)]$ and concave over $[u/(\beta\ell),1]$. [^13] Kolotilin et al. (2022) describe such a $\phi$ function as “S-shaped.” They show that if $\phi$ is continuously differentiable and S-shaped, there exists a solution to (LP-$u$) that is an upper-censorship: perfectly segment types below a cutoff and pool all the types above the cutoff. However, their results cannot be directly applied here as the $\phi(x,u)$ is not differentiable at $x=\frac{u}{\beta\ell}$. Additionally, Proposition 5 goes further than an upper censorship; it claims that a binary segmentation is sufficient to obtain constrained efficiency.
@@
Next, define the cutoff $\theta^{c}_{u}\coloneqq\min\{\theta\in[\underline{\theta}_{u},1]:G_{u}(\theta)\geq 0\}$, which is well-defined because $G_{u}$ is continuous and strictly increasing, and $\{\theta\in[\underline{\theta}_{u},1]:G_{u}(\theta)\geq 0\}$ is non-empty. [^15] For example, $01-\overline{X}(1)=$ 0, so $G_{u}(1)=g(\ell/u)>0$ for each $u\in(0,\beta\ell)$. The inequality follows because $g$ is a strictly increasing function with $g(1/\beta)=0$.
@@
where the inequality follows from the concavity of $g$, and the last equality follows because $g(1/\beta)=0$. Since $\theta^{c}_{u}$ is the smallest type in $[\underline{\theta}_{u},1]$ that satisfies $G_{u}(\theta)\geq 0$, we have $u/(\beta\ell)\geq\theta_{u}^{c}$.
@@
[…]

### v3

@@ 4. $(iv)$
where the first equality follows from the definition of $\sigma^{PS}$ and the characterization of $\tau^{FB}_{\sigma^{PS}}$ in Proposition 1, the second equality follows by the assumption on $\lambda$ given in Proposition 4 and the fact that $f$ is the inverse of the mapping $t\mapsto 1/m^{\prime}(t)$, and the last equality follows by the construction of $\eta_{\sigma^{PS}}$ as given in Proposition 1. Thus, the pair $(\tau^{FB}_{\sigma^{PS}},\lambda({\eta_{\sigma^{PS}}}/{\beta})\cdot\eta_{\sigma^{PS}})$ also satisfies (4) of Definition 1.
We can therefore conclude that $(\tau^{FB}_{\sigma^{PS}},\lambda({\eta_{\sigma^{PS}}}/{\beta})\cdot\eta_{\sigma^{PS}})$ is a search equilibrium of $\sigma^{PS}$. However, $(\tau^{*}_{\sigma^{PS}},u^{*}_{\sigma^{PS}})$ is the essentially-unique search equilibrium of $\sigma^{PS}$. Thus, $\tau^{*}_{\sigma^{PS}}=\tau^{FB}_{\sigma^{PS}}$ for $\sigma_{\mu}^{PS}$-almost everywhere, and $u^{*}_{\sigma^{PS}}=\lambda({\eta_{\sigma^{PS}}}/{\beta})\cdot\eta_{\sigma^{PS}}$. Since $(\sigma^{PS},\tau^{FB}_{\sigma^{PS}})$ solves (FB), the almost-everywhere equivalence between $\tau^{*}_{\sigma^{PS}}$ and $\tau^{FB}_{\sigma^{PS}}$ implies that $(\sigma^{PS},\tau^{*}_{\sigma^{PS}})$ also solves (FB), giving us the desired result. ∎
@@
$\max_{\widehat{H}\in\operatorname*{MPC}(F)}\hskip 5.0ptk\int_{\Theta}\phi(x,u)\hskip 1.00006ptd\widehat{H}(x).$ (LP-$u$)
Let V(u){V}(u) be the value function corresponding to (LP-$u$), and let $\mathcal{H}(u)$ be the set of maximizers. Since the objective is a continuous function and $\operatorname*{MPC}(F)$ is compact, $V(u)$ is continuous in $u$, and $\mathcal{H}(u)$ is non-empty and compact-valued for each $u>0$. Importantly, by construction, $H$ satisfies (5) if and only if $H\in\mathcal{H}(u_{H}^{*})$.
@@
$\underbrace{\ell\cdot k\int_{\Theta}m(\tau_{H}^{*}(x))xdH(x)}_{=u_{H}^{*}}\geq\underbrace{\ell\cdot k\int_{\Theta}m(\tau_{\widehat{H}}^{*}(x))xd\widehat{H}(x)}_{=u_{\widehat{H}}^{*}},$
@@
First, suppose the mapping $t\mapsto t/m(t)$ is concave, and therefore, $g$ is convex. In this case, $\phi(x,u)=g({\ell\cdot x}/{u})\mathbbm{1}_{[\beta\ell x>u]}$ is convex in $x$ for all $u\in(0,\beta\ell)$. [^13] Recall that $\lim_{y\to 1/\beta}g(y)=0$. As a result, for all $H^{\prime},H^{\prime\prime}\in\operatorname*{MPC}(F)$ where $H^{\prime}$ is a mean-preserving contraction of $H^{\prime\prime}$, we have
$k\int_{\Theta}\phi(x,u)dH^{\prime}(x)\leq k\int_{\Theta}\phi(x,u)dH^{\prime\prime}(x).$
@@
Second, suppose the mapping $t\mapsto t/m(t)$ is convex, and therefore, $g$ is concave. In this case, $\phi(x,u)=g({\ell\cdot x}/{u})\mathbbm{1}_{[\ell\beta x>u]}$ is piece-wise concave in $x$ for all $u\in(0,\beta\ell)$. Specifically, it is constant over $0,u/(\beta\ell)]$ and concave over $u/(\beta\ell),1]$. [^14] Kolotilin et al. (2022) describe such a $\phi$ function as “S-shaped.” They show that if $\phi$ is continuously differentiable and S-shaped, there exists a solution to (LP-$u$) that is an upper-censorship: perfectly segment types below a cutoff and pool all the types above the cutoff. However, their results cannot be directly applied here as the $\phi(x,u)$ is not differentiable at $x=\frac{u}{\beta\ell}$. Additionally, Proposition 5 goes further than an upper censorship; it claims that a binary segmentation is sufficient to obtain constrained efficiency.
@@
Next, define the cutoff $\theta^{c}_{u}\coloneqq\min\{\theta\in[\underline{\theta}_{u},1]:G_{u}(\theta)\geq 0\}$, which is well-defined because $G_{u}$ is continuous and strictly increasing, and $\{\theta\in[\underline{\theta}_{u},1]:G_{u}(\theta)\geq 0\}$ is non-empty. [^16] For example, $1-\overline{X}(1)=0$, so $G_{u}(1)=g(\ell/u)>0$ for each $u\in(0,\beta\ell)$. The inequality follows because $g$ is a strictly increasing function with $g(1/\beta)=0$.
@@
where the inequality follows from the concavity of $g$, and the last equality follows because $g(1/\beta)=0$. Since $\theta^{c}_{u}$ is the smallest type in $\underline{\theta}_{u},1]$ that satisfies $G_{u}(\theta)\geq 0$, we have $u/(\beta\ell)\geq\theta_{u}^{c}$.
@@ 3. $()$
$\mathbb{E}_{F}[p_{u}(x)]=$ $\int^{1}_{\theta^{c}_{u}}\left[g\left(\frac{\ell\cdot\overline{X}(\theta^{c}_{u})}{u}\right)+g^{\prime}\left(\frac{\ell\cdot\overline{X}(\theta^{c}_{u})}{u}\right)\frac{\ell}{u}\Big{(}x-\overline{X}(\theta^{c}_{u})\Big{)}\right]dF(x)$
$=$ $p_{u}\big{(}\overline{X}(\theta^{c}_{u})\big{)}(1-F(\theta^{c}_{u}))+g^{\prime}\left(\frac{\ell\cdot\overline{X}(\theta^{c}_{u})}{u}\right)\frac{\ell}{u}\underbrace{\int^{1}_{\theta^{c}_{u}}\Big{(}x-\overline{X}(\theta^{c}_{u})\Big{)}dF(x)}_{=0}$